OpenStudy (anonymous):

How would you find the horizontal asymptote of f(x)=(x^2-16)/(x+3) Thanks!!

OpenStudy (anonymous):

differentiate and put \[f \prime(x)=0\] find value of x and then f(x) y=f(x) will be the answer

OpenStudy (anonymous):

Okay, so then I honestly still have no idea what to actually write as my answer

OpenStudy (anonymous):

y=2or -8 two of them would be there

OpenStudy (anonymous):

okay, thanks!!

OpenStudy (anonymous):

medal please

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I was getting there :P

OpenStudy (anonymous):

there is a small mistake that i made the answer is y= -48/5 or -12/5

OpenStudy (anonymous):

@annas My question is Solve 3^(x-4) = 7^(x+9) Thanks!

OpenStudy (anonymous):

ok.... np

OpenStudy (anonymous):

\[3^{x-4} = 7^{x+9}\]

OpenStudy (anonymous):

(x-4)log3=(x+9)log7 (x-4)(0.4771)=(x+9)(0.8450) 0.4771x-1.908=0.8450x+7.605 -1.908-7.605=0.3679x

OpenStudy (anonymous):

The answers are A x ≈ –2.59 B x ≈ –25.86 C x ≈ –0.39 D x ≈ –0.04

OpenStudy (anonymous):

-9.513=0.3679x x=-25.857

OpenStudy (anonymous):

its B

OpenStudy (anonymous):

Thank you so much!!!!! How else can I give you a medal?

OpenStudy (anonymous):

no need of medals i dont work for medals i love helping boy :) and welcome

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

no need of that either boy ...